# RaffeSDG: Random Frequency Filtering enabled Single-source Domain Generalization for Medical Image Segmentation

### v2

@@ IV-E1 Data volume impact
IV-E2 Ablation against segmentor modules
@@
TABLE VI: Comparisons of OCT and ultrasound segmentation. The top three results are indicated in red, magenta, and blue.
@@
SAN-SAW [16] 0.860 0.613 0.771 0.748 0.735 0.825 0.750 0.770
SLAug [13] 0.483 0.316 0.458 0.419 0.745 0.830 0.765 0.780
GIN-IPA [14] 0.952 0.569 0.703 0.743 0.715 0.834 0.781 0.777
FreeSDG [15] 0.885 0.565 0.760 0.736 0.750 0.826 0.759 0.778
RaffeSDG(ours) 0.901 0.579 0.778 0.753 0.766 0.847 0.793 0.802
@@
As observed in Fig. 10, the introduction of data augmentation significantly enhances out-of-domain segmentation by the backbone module. Then, the self-supervision based on Gaussian filters imposes additional constraints across augmented samples, enabling the learning of domain-invariant representations within the segmontor. Afterward, the attention blocks efficiently incorporate the representations learned from the pretext tasks, resulting in a significant enhancement in out-of-domain inference capabilities of the segmentor.
IV-F Versatility for various tissues and imaging modalities
Comparisons of out-of-domain inference for OCT and ultrasound image segmentation were also performed to validate the effectiveness of RaffeSDG. Visualized and quantitative summaries of these comparisons are respectively presented in Fig. 9 and Table VI.
Considering that SAN-SAW [16] and GIN-IPA [14] employ specific designs for multi-category labels, it is understandable that they achieve remarkable performance in the retina layer segmentation of OCT, as expressed in Table VI. However, the nature of OCT and ultrasound data poses limitations for RaffeSDG. Unlike colorful fundus data obtained through visible light, OCT and ultrasound utilize wave propagation characteristics to capture images of internal structures, resulting in an inherently grayscale appearance. Consequently, OCT and ultrasound do not lend themselves well to the advantages of RaffeSDG, as illustrated in Fig. 9. Fortunately, despite these constraints, RaffeSDG still manages to achieve decent results, highlighting its versatility in medical image segmentation.
V Discussions

### v3

@@ 4.5.1 Data volume impact
4.5.2 Ablation against segmentor modules
@@
Table 6: Comparisons of OCT and ultrasound segmentation. The top three results are indicated in red, magenta, and blue.
@@
SLAug [su2022rethinking] 0.961 0.831 0.939 0.910 0.613 0.639 0.638 0.630
GIN-IPA [ouyang2022causality] 0.983 0.955 0.941 0.960 0.632 0.636 0.626 0.631
CCDR [liang2024single] 0.980 0.939 0.951 0.956 0.619 0.638 0.640 0.632
FreeSDG [li2023frequency] 0.978 0.947 0.938 0.955 0.648 0.614 0.624 0.629
RaffeSDG(ours) 0.981 0.953 0.960 0.964 0.644 0.663 0.658 0.655
@@
As observed in Fig. 10, the introduction of data augmentation significantly enhances out-of-domain segmentation by the backbone module. Then, the self-supervision based on Gaussian filters imposes additional constraints across augmented samples, enabling the learning of domain-invariant representations within the segmentor. Afterward, the attention blocks efficiently incorporate the representations learned from the pretext tasks, resulting in a significant enhancement in out-of-domain inference capabilities of the segmentor.
4.6 Versatility for various tissues and imaging modalities
Comparisons of out-of-domain inference for OCT and ultrasound image segmentation were also performed to validate the effectiveness of RaffeSDG. Visualized and quantitative summaries of these comparisons are respectively presented in Fig. 9 and Table 6.
Considering that CCDR [liang2024single] and GIN-IPA [ouyang2022causality] employ specific designs for multi-category labels, it is understandable that they achieve remarkable performance in the retina layer segmentation of OCT, as expressed in Table 6. However, the nature of OCT and ultrasound data poses limitations for RaffeSDG. Unlike colorful fundus data obtained through visible light, OCT and ultrasound utilize wave propagation characteristics to capture images of internal structures, resulting in an inherently grayscale appearance. Consequently, OCT and ultrasound do not lend themselves well to the advantages of RaffeSDG, as illustrated in Fig. 9. Fortunately, despite these constraints, RaffeSDG still manages to achieve decent results, highlighting its versatility in medical image segmentation.
5 Discussions
@@ 6 Conclusions
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (62401246), Guangdong Basic and Applied Basic Research Foundation (2514050003407) and Shenzhen Natural Science Fund (JCYJ20240813095112017).
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, we used DeepSeek-R1 in order to optimize textual expression and clarity (including fluency, grammar, word choice consistency, and the readability of the abstract and title). After using this tool/service, we reviewed and edited the content as needed and take full responsibility for the content of the published article. The tool was not used for study design, data collection, data analysis, or the generation of results, figures, or references.